OpenStudy (anonymous):

I am trying to help a neighbor solving this, he's a 9th grader, but I don't remember algebra, can you help me with this equation. Solve each proportion p/4 = 7/8

OpenStudy (anonymous):

I think p=3.5

OpenStudy (anonymous):

4 is half of 8, and half of 7 is 3.5

OpenStudy (anonymous):

4*7=28/8=3.5